# Towards Robust Multimodal Sentiment Analysis with Incomplete Data

### v1

@@ Abstract
The field of Multimodal Sentiment Analysis (MSA) has recently witnessed an emerging direction seeking to tackle the issue of data incompleteness. Recognizing that the language modality typically contains dense sentiment information, we consider it as the dominant modality and present an innovative Language-dominated Noise-resistant Learning Network (LNLN) to achieve robust MSA. The proposed LNLN features a dominant modality correction (DMC) module and dominant modality based multimodal learning (DMML) module, which enhances the model’s robustness across various noise scenarios by ensuring the quality of dominant modality representations. Aside from the methodical design, we perform comprehensive experiments under random data missing scenarios, utilizing diverse and meaningful settings on several popular datasets (e.g., MOSI, MOSEI, and SIMS), providing additional uniformity, transparency, and fairness compared to existing evaluations in the literature. Empirically, LNLN consistently outperforms existing baselines, demonstrating superior performance across these challenging and extensive evaluation metrics. Our code will be available at: https://github.com/Haoyu-ha/LNLN.
@@ 1 Introduction
Changing trends in recent research have taken a turn towards modeling data in natural scenarios from laboratory conditions (Tsai et al., 2019; Hazarika et al., 2020; Yu et al., 2021; Zhang et al., 2023). The shift has created a wider application space in the real-world for MSA, though concerns arise due to problems like sensor failures and problems with Automatic Speech Recognition (ASR), leading to inconsistencies such as incomplete data in real-world deployment.
@@
Addressing this gap, our paper aims to offer a comprehensive evaluation on three widely-used datasets, namely MOSI (Zadeh et al., 2016), MOSEI (Zadeh et al., 2018) and SIMS (Yu et al., 2020) datasets. We introduce random data missing instances and subsequently compare the performance of existing methods on these datasets. This endeavor seeks to provide an all-encompassing outlook for evaluating the effectiveness and robustness of various methods in the face of incomplete data, thereby sparking new insights in the field.
Additionally, inspired by the previous work ALMT (Zhang et al., 2023), we hypothesize that model robustness improves when the integrity of the dominant modality is preserved despite varying noise levels. Therefore, we introduce a novel model, namely Language-dominated Noise-resistant Learning Network (LNLN), to enhance MSA’s robustness over incomplete data. LNLN aims to augment the integrity of the language modality’s features, regarded as the dominant modality due to its richer sentiment cues, with the support of other auxiliary modalities. The LNLN’s robustness against varying levels of data incompleteness is achieved through a dominant modality correction (DMC) module for dominant modality construction, a dominant modality based multimodal learning (DMML) module for multimodal fusion and classification, and a reconstructor for reconstructing missing information to shield the dominate modality from noise interference. This approach ensures a high-quality dominant modality feature, which significantly bolsters the robustness of LNLN under diverse noise conditions. Consequently, extensive experimental results demonstrate the LNLN’s superior performance across these challenging and extensive evaluation metrics.
In summary, this paper conducts a comprehensive evaluation of existing advanced MSA methods with missing data. This analysis highlights the strengths and weaknesses of various approaches when contending with incomplete data. We believe, it can improve the understanding of different MSA methods’ performance under complex real-world scenarios, thereby informing technology’s future trajectory. Our proposed LNLN also offers valuable insight and guidelines for thriving in this research space.
@@ 2.2 Robust Representation Learning in MSA
Context-based MSA and Noise-aware MSA differ in their approaches to robust representation learning. In Context-based MSA, robust representation learning typically relies on modeling intra- and inter-modality relationships. For instance, Hazarika et al. (2020) and Yang et al. (2022) apply feature disentanglement to each modality, modeling multimodal representations from multiple feature subspaces and perspectives. Yu et al. (2021) and Liang et al. (2020) explore self-supervised learning and semi-supervised learning to enhance multimodal representations, respectively. Tsai et al. (2019) and Rahman et al. (2020) introduce Transformer to learn the long dependencies of modalities. Zhang et al. (2023) devise a language-guided learning mechanism that uses modalities with more intensive sentiment cues to guide the learning of other modalities.
In contrast, Noise-aware MSA focuses more on perceiving and eliminating the noise present in the data. For example, For example, Mittal et al. (2020) design a modality check module based on metric learning and Canonical Correlation Analysis (CCA) to identify the modality with greater noise. Yuan et al. (2021) design a feature reconstruction network to predict the location of missing information in sequences and reconstruct it. Yuan et al. (2024) introduce adversarial learning (Goodfellow et al., 2014) to perceive and generate cleaner representations.
In our work, the LNLN belongs to the Noise-aware MSA category. Inspired by Zhang et al. (2023), we explore the capability of language-guided mechanisms in resisting noise and aim to provide new perspectives for the study of MSA in noisy scenarios.
@@ 3.1 Overview
The overall pipeline for the proposed Language-dominated Noise-resistant Learning Network (LNLN) in robust multimodal sentiment analysis is illustrated in Figure 1. As depicted, a crucial initial step involves forming a multimodal input with random data missing, which sets the stage for LNLN training. Once the input is prepared, LNLN first utilizes an embedding layer to standardize the dimension of each modality, ensuring uniformity. Recognizing that language is the dominant modality in MSA (Zhang et al., 2023), a specially designed Dominant Modality Correction (DMC) module employs adversarial learning and a dynamic weighted enhancement strategy to mitigate noise impacts. This module first enhances the quality of the dominant feature computed from the language modality and then integrates them with the auxiliary modalities (visual and audio) in the dominant modality based multimodal learning (DMML) module for effective multimodal fusion and classfication. This process significantly bolsters LNLN’s robustness against various noise levels. Moreover, to refine the network’s capability for fine-grained sentiment analysis, a simple reconstructor is implemented to reconstruct missing data, further enhancing the system’s robustness.
@@ 3.2 Input Construction and Multimodal Input
Given the challenges of random data missing, we have constructed data sets that simulate these conditions based on MOSI, MOSEI and SIMS datasets.
Random Data Missing. Following previous protocols (Yuan et al., 2021), for each modality in each sample, we randomly erased changing proportions of information (from 0% to 100%). Specifically, for visual and audio modalities, we fill the erased information with zeros. For language modality, we fill the erased information with [UNK], which indicates the unknown word in BERT (Kenton and Toutanova, 2019).
Multimdoal Input. For each sample in the dataset, we incorporate data from three modalities: language, audio, and visual data. Consistent with previous works (Zhang et al., 2023), each modality is processed using widely-used tools: language data is encoded using BERT (Kenton and Toutanova, 2019), audio features are extracted through Librosa (McFee et al., 2015), and visual features are obtained using OpenFace (Baltrusaitis et al., 2018). These pre-processed inputs are represented as sequences, denoted by $U^{0}_{m}\in\mathbb{R}^{T_{m}\times d_{m}}$, where $m\in\{l,v,a\}$ represents the modality type ($l$ for language, $v$ for visual, $a$ for audio), $T_{m}$ indicates the sequence length, and $d_{m}$ refers to the dimension of each modality’s vector. With obtained $U^{0}_{m}$, we apply random data missing to $U^{0}_{m}$, thus forming the noise corrupted multimodal input $U^{1}_{m}\in\mathbb{R}^{T_{m}\times d_{m}}$.
@@ 3.3 Dominant Modality based Multimodal Learning
Inspired by previous work ALMT, we hypothesize that model robustness improves when the integrity of the dominant modality is preserved despite varying noise levels. We improve ALMT based on the designed DMC module and Reconstructor, thus implementing dominant modality based DMML module for sentiment analysis under random data missing scenarios. Here, we mainly introduce the parts that differ from ALMT. Further details available in Zhang et al. (2023).
@@
$H_{\text{hyper}}^{i}=H_{\text{hyper }}^{i-1}+\operatorname{MHA}(H_{d}^{i},H_{a}^{1})+\operatorname{MHA}(H_{d}^{i},H_{v}^{1}),$ (3)
where $\operatorname{MHA}(\cdot)$ represents multi-head attention, $H_{\text{hyper}}^{i}\in\mathbb{R}^{T\times d}$ is the hyper-modality feature. Note that the feature $H_{\text{hyper}}^{0}\in\mathbb{R}^{T\times d}$ is a random initialized vector.
@@ 3.4 Dominant Modality Correction
Completeness Check. We apply an encoder $\operatorname{E}_{cc}$ that consists of a Transformer encoder with the depth of two layers and a classifier for completeness check. For example, if the missing rate of dominate modality is 0.3, the label of completeness is 0.7. This completeness prediction $w$ can be obtained as follows:
@@ 3.6 Overall Learning Objectives
where $\alpha$, $\beta$, $\gamma$ and $\delta$ are hyper-parameters. On MOSI and MOSEI datasets, we empirically set them to 0.9, 0.8, 0.1, and 1.0, respectively. On SIMS dataset, we empirically set them to 0.9, 0.6, 0.1, and 1.0, respectively.
@@ 4 Experiments and Analysis
In this section, we provide a comprehensive and fair comparison between the proposed LNLN and previous state-of-the-art MSA methods on MOSI (Zadeh et al., 2016), MOSEI (Zadeh et al., 2018) and SIMS (Yu et al., 2020) datasets. 1) For each method, we select three random seeds for training and evaluation. 2) We report the performance of the model on a variety of evaluation metrics, which is much more comprehensive compared to some of the previous work.
@@ 4.1 Datasets
MOSEI. The dataset consists of 22,856 video clips sourced from YouTube. The sample is divided into 16,326 clips for training, 1,871 for validation, and 4,659 for testing. Each clip is carefully labeled with a sentiment score, ranging from -3, representing the strongly negative sentiment, to 3, denoting the strongly positive sentiment.
@@ 4.2 Evaluation Settings and Criteria
For evaluation criteria, we report the binary classification accuracy (Acc-2), the F1 score associated with Acc-2, and the mean absolute error (MAE). For Acc-2, we calculated accuracy and F1 in two ways: negative/non-negative and negative/positive on the MOSI and MOSEI datasets, respectively. Additionally, we provide the three-class accuracy (Acc-3), the seven-class accuracy (Acc-7), and the correlation of the model’s prediction with humans (Corr) on the MOSI dataset. For the SIMS dataset, we report Acc-3, the five-class accuracy (Acc-5), and Corr.
4.3 Robustness Comparison
Tables 1 and 2 show the robustness evaluation results on the MOSI, MOSEI, and SIMS datasets. As shown in Table 1, LNLN achieves state-of-the-art performance on most metrics. For example, on the MOSI dataset, LNLN achieved a relative improvement of 9.46% on Acc-7 compared to the sub-optimal result obtained by MMIM, demonstrating the robustness of LNLN in the face of different noise effects. However, on the MOSEI dataset, LNLN achieves only sub-optimal performance on metrics such as Acc-7 and Acc-5. After analyzing the data distribution (see Appendix B.3) and the confusion matrix (see Appendix B.5), we believe that this is due to the fact that LNLN does not overly bias the neutral category, which has a disproportionate number of samples in noisy scenarios.
As shown in Table 2, LNLN achieves a remarkable improvement in F1 on the SIMS dataset, with a relative improvement of 4.23% on F1 compared to the sub-optimal result obtained by ALMT. Similar to CENET’s performance on the MOSEI dataset, TETFN on the SIMS dataset has a strong tendency to predict inputs as weakly negative categories with high sample sizes, resulting in seemingly better performance on some metrics.
Additionally, as shown in Figure 2, we present the performance curves of several advanced methods under varying missing rates. The results demonstrate that the proposed LNLN consistently outperforms others across most scenarios, showing its robustness under different missing rates.
Table 1: Robustness comparison of the overall performance on MOSI and MOSEI datasets. Note: The smaller MAE indicate the better performance.
Method MOSI MOSEI
Acc-7 Acc-5 Acc-2 F1 MAE Corr Acc-7 Acc-5 Acc-2 F1 MAE Corr
Table 2: Robustness comparison of the overall performance on SIMS dataset. Note: The smaller MAE indicate the better performance.
Method Acc-5 Acc-3 Acc-2 F1 MAE Corr
MISA 31.53 56.87 72.71 66.30 0.539 0.348
Self-MM 32.28 56.75 72.81 68.43 0.508 0.376
MMIM 31.81 52.76 69.86 66.21 0.544 0.339
CENET 22.29 53.17 68.13 57.90 0.589 0.107
TETFN 33.42 56.91 73.58 68.67 0.505 0.387
TFR-Net 26.52 52.89 68.13 58.70 0.661 0.169
ALMT 34.16 56.47 71.85 76.21 0.509 0.372
LNLN 34.64 57.14 72.73 79.43 0.514 0.397
Figure 2: Performance curves of various missing rates. (a), (b) and (c) is the F1 curves on MOSI, MOSEI and SIMS, respectively. (d), (e) and (f) is the MAE curves on MOSI, MOSEI and SIMS, respectively. Note: The smaller MAE indicate the better performance.
4.4 Effects of Different Components
To verify the role of different modules, we present the ablation results after subtracting each component on the MOSI dataset separately. As shown in Table 3, a few metrics show an upward trend in performance when certain modules are removed. On one hand, we believe this is due to randomness in noisy scenarios. On the other hand, the same hyper-parameters cannot make the model perform optimally across all metrics.
Moreover, we can see that the performance of LNLN decreases relatively slightly when the DMC and reconstructor are removed individually. However, there is a significant drop in performance when both are removed, e.g., 4.72% decrease in Acc-7 and 5.87% decrease in Acc-5. Performance drops further after replacing DMML with concatenation fusion, which also proves that DMML plays an important role in MSA. Additionally, after removing the noisy data used for training, the performance of LNLN also shows a significant decrease. This observation indicates that LNLN needs to learn to be aware of noise with the help of noisy data.
Overall, both the presence of noisy data and well-designed components are crucial for robust MSA in scenarios with random missing data.
Table 3: Effects of different components. Note: The smaller MAE indicate the better performance.
Method MOSI SIMS
Acc-7 Acc-5 Acc-2 F1 MAE Corr Acc-5 Acc-3 Acc-2 F1 MAE Corr
4.5 Effects of Different Regularization
As shown in Table 4, we present the ablation results after subtracting each regularization on the MOSI dataset separately. Obviously, after each regularization is removed, LNLN shows a decrease in performance on most metrics. This demonstrates that regularization contributes to the learning of the model. Similar to the results in Table 3, there are a few metrics that improve after the removal of regularization. We believe this is mainly due to the difficulty of optimizing a single set of hyper-parameters to achieve the best performance across all metrics.
Method MOSI SIMS
Acc-7 Acc-5 Acc-2 F1 MAE Corr Acc-5 Acc-3 Acc-2 F1 MAE Corr
4.6 Effects of Different Modalities
To verify the role of different modalities, we present the ablation results after subtracting each modality from the MOSI dataset. As shown in Table 5, all modalities contribute to the final performance. Notably, when the language modality is removed, the performance of LNLN shows a significant decrease, while removing the visual and audio modalities results in a smaller decrease. This indicates that the language modality plays a crucial role in the MSA dataset.
Moreover, we observe that LNLN converges on most metrics even when the visual and audio modalities are removed. This is due to the model’s lazy behavior, where LNLN tends to fix its predictions to a certain categories. More details can be found in Appendix B.3 and B.5.
Table 5: Effects of different modalities. Note: The smaller MAE indicate the better performance.
Method MOSI
Acc-7 Acc-5 Acc-2 F1 MAE Corr
LNLN 34.26 38.27 72.55 / 70.94 72.73 / 71.25 1.046 0.527
w/o language 16.30 16.31 54.88 / 55.79 60.33 / 63.10 1.399 0.033
w/o audio 33.51 38.10 72.31 / 70.54 73.67 / 71.10 1.064 0.535
w/o vision 33.53 38.15 72.32 / 70.58 73.68 / 71.14 1.064 0.536
4.7 Case Study
As shown in Figure 3, we visualize several successful and failed predictions made by LNLN and ALMT from MOSI dataset for case study. It shows that LNLN can perceives sentiment cues in challenging samples, which demonstrates its ability to capture sentiment cues in noisy scenario. However, for inputs with high missing rates (e.g., the third example in the figure), both LNLN and ALMT fail to make correct predictions. We believe this is due to the high loss of valid information in the multimodal input, making accurate predictions difficult.
Figure 3: Visualization of successful and failure predictions.
4.8 Discussion
In this work, we conduct a comprehensive evaluation of current advanced methods, the results of which have triggered some thoughts. Specifically, we find that all methods fail to accurately predict the inputs in high missing rate scenarios because most of them are designed specifically for complete data (see Appendix B.5 for more details). However, some methods show relatively more robust performance in low missing rate scenarios and some specific scenarios (see Appendix B.7), such as Self-MM, TETFN, and ALMT. We believe that some of the ideas in these methods are useful for future research on more robust MSA for incomplete data.
Additionally, we explore the performance of these methods in modality missing scenarios, which is a special case of random data missing where the partial modality missing rate $r=1.0$ (see Appendix B.7 for more details). We find that ALMT and LNLN demonstrate advantages over other methods in terms of F1 metrics when the language modality is completely lost.
5 Conclusion and Future Work
In this paper, a novel Language-dominated Noise-resistant Learning Network (LNLN) is proposed to better learn sentiment cues for robust MSA. Due to the collaboration between Dominate Modality Correction (DMC) module and Dominant Modality based Multimodal Learning (DMML) module, and Reconstructor, LNLN can dynamically enhancing the dominant modality to achieve superior performance in data missing scenarios with varying levels. Extensive evaluation shows that none of the existing methods can effectively model the data with high missing rates. The related analyses can provide suggestions for other researchers to better handle robust MSA. In the future, we will focus on improving the generalization of the model to handle different types of scenes and varying intensities of noise effectively.
Limitations
We believe there are several limitations to the LNLN: 1) The LNLN achieves well performance in data missing scenarios, but not always better than all other methods in the modality missing scenarios, which demonstrates the lack of multi-scene generalization capability of the LNLN. 2) The data in real-world scenarios is much more complex. In addition to the presence of missing data, other factors need to be considered, such as diverse cultural contexts, varying user behavior patterns, and the influence of platform-specific features on the data. These factors can introduce additional noise and variability, which may require further model adaptation and tuning to handle effectively. 3) Tuning the hyperparameters, particularly those related to the loss functions, can be challenging and may require more sophisticated methods to achieve optimal performance.
Appendix A Implementation Details

### v2

@@ Abstract
The field of Multimodal Sentiment Analysis (MSA) has recently witnessed an emerging direction seeking to tackle the issue of data incompleteness. Recognizing that the language modality typically contains dense sentiment information, we consider it as the dominant modality and present an innovative Language-dominated Noise-resistant Learning Network (LNLN) to achieve robust MSA. The proposed LNLN features a dominant modality correction (DMC) module and dominant modality based multimodal learning (DMML) module, which enhances the model’s robustness across various noise scenarios by ensuring the quality of dominant modality representations. Aside from the methodical design, we perform comprehensive experiments under random data missing scenarios, utilizing diverse and meaningful settings on several popular datasets (e.g., MOSI, MOSEI, and SIMS), providing additional uniformity, transparency, and fairness compared to existing evaluations in the literature. Empirically, LNLN consistently outperforms existing baselines, demonstrating superior performance across these challenging and extensive evaluation metrics. [^1] The code is available at: https://github.com/Haoyu-ha/LNLN
@@ 1 Introduction
Changing trends in recent research has taken a turn towards modeling data in natural scenarios from laboratory conditions (Tsai et al., 2019; Hazarika et al., 2020; Yu et al., 2021; Zhang et al., 2023). The shift has created a wider application space in the real-world for MSA, though concerns arise due to problems like sensor failures and problems with Automatic Speech Recognition (ASR), leading to inconsistencies such as incomplete data in real-world deployment.
@@
Addressing this gap, our paper aims to offer a comprehensive evaluation on three widely-used datasets, namely MOSI (Zadeh et al., 2016), MOSEI (Zadeh et al., 2018) and SIMS (Yu et al., 2020) datasets. We introduce random data missing instances and subsequently compare the performance of existing methods on these datasets. This endeavor seeks to provide an all-encompassing outlook for evaluating the effectiveness and robustness of various methods in the face of incomplete data, thereby sparking new insights in the field. Additionally, inspired by the previous work ALMT (Zhang et al., 2023), we hypothesize that model robustness improves when the integrity of the dominant modality is preserved despite varying noise levels. Therefore, we introduce a novel model, namely Language-dominated Noise-resistant Learning Network (LNLN), to enhance MSA’s robustness over incomplete data. LNLN aims to augment the integrity of the language modality’s features, regarded as the dominant modality due to its richer sentiment cues, with the support of other auxiliary modalities. The LNLN’s robustness against varying levels of data incompleteness is achieved through a dominant modality correction (DMC) module for dominant modality construction, a dominant modality based multimodal learning (DMML) module for multimodal fusion and classification, and a reconstructor for reconstructing missing information to shield the dominate modality from noise interference. This approach ensures a high-quality dominant modality feature, which significantly bolsters the robustness of LNLN under diverse noise conditions. Consequently, extensive experimental results demonstrate the LNLN’s superior performance across these challenging and evaluation metrics.
In summary, this paper conducts a comprehensive evaluation of existing advanced MSA methods. This analysis highlights the strengths and weaknesses of various methods when contending with incomplete data. We believe, it can improve the understanding of different MSA methods’ performance under complex real-world scenarios, thereby informing technology’s future trajectory. Our proposed LNLN also offers valuable insight and guidelines for thriving in this research space.
@@ 2.2 Robust Representation Learning in MSA
Context-based MSA and Noise-aware MSA differ in their approaches to robust representation learning. In Context-based MSA, robust representation learning typically relies on modeling intra- and inter-modality relationships. For instance, Hazarika et al. (2020) and Yang et al. (2022) apply feature disentanglement to each modality, modeling multimodal representations from multiple feature subspaces and perspectives. Yu et al. (2021) and Liang et al. (2020) explore self-supervised learning and semi-supervised learning to enhance multimodal representations, respectively. Tsai et al. (2019) and Rahman et al. (2020) introduce Transformer to learn the long dependencies of modalities. Zhang et al. (2023) devise a language-guided learning mechanism that uses modalities with more intensive sentiment cues to guide the learning of other modalities. In contrast, Noise-aware MSA focuses more on perceiving and eliminating the noise present in the data. For example, Mittal et al. (2020) design a modality check module based on metric learning and Canonical Correlation Analysis (CCA) to identify the modality with greater noise. Yuan et al. (2021) design a feature reconstruction network to predict the location of missing information in sequences and reconstruct it. Yuan et al. (2024) introduce adversarial learning (Goodfellow et al., 2014) to perceive and generate cleaner representations.
In this work, the LNLN belongs to the noise-aware MSA category. Inspired by Zhang et al. (2023), we explore the capability of language-guided mechanisms in resisting noise and aim to provide new perspectives for the study of MSA in noisy scenarios.
@@ 3.1 Overview
The overall pipeline for the proposed Language-dominated Noise-resistant Learning Network (LNLN) in robust multimodal sentiment analysis is illustrated in Figure 1. As depicted, a crucial initial step involves forming a multimodal input with random data missing, which sets the stage for LNLN training. Once the input is prepared, LNLN first utilizes an embedding layer to standardize the dimension of each modality, ensuring uniformity. Recognizing that language is the dominant modality in MSA (Zhang et al., 2023), a specially designed Dominant Modality Correction (DMC) module employs adversarial learning and a dynamic weighted enhancement strategy to mitigate noise impacts. This module first enhances the quality of the dominant feature computed from the language modality and then integrates them with the auxiliary modalities (visual and audio) in the dominant modality based multimodal learning (DMML) module for effective multimodal fusion and classification. This process significantly bolsters LNLN’s robustness against various noise levels. Moreover, to refine the network’s capability for fine-grained sentiment analysis, a simple reconstructor is implemented to reconstruct missing data, further enhancing the system’s robustness.
@@ 3.2 Input Construction and Multimodal Input
Given the challenges of random data missing, we have constructed data sets that simulate these conditions based on MOSI, MOSEI, and SIMS datasets.
Random Data Missing. Following the previous method (Yuan et al., 2021), for each modality, we randomly erased changing proportions of information (from 0% to 100%). Specifically, for visual and audio modalities, we fill the erased information with zeros. For language modality, we fill the erased information with [UNK] which indicates the unknown word in BERT (Kenton and Toutanova, 2019).
Multimdoal Input. For each sample in the dataset, we incorporate data from three modalities: language, audio, and visual data. Consistent with previous works (Zhang et al., 2023), each modality is processed using widely-used tools: language data is encoded using BERT (Kenton and Toutanova, 2019), audio features are extracted through Librosa (McFee et al., 2015), and visual features are obtained using OpenFace (Baltrusaitis et al., 2018). These pre-processed inputs are represented as sequences, denoted by $U^{0}_{m}\in\mathbb{R}^{T_{m}\times d_{m}}$, where $m\in\{l,v,a\}$ represents the modality type ($l$ for language, $v$ for visual, $a$ for audio), $T_{m}$ indicates the sequence length and $d_{m}$ refers to the dimension of each modality’s vector. With obtained $U^{0}_{m}$, we apply random data missing to $U^{0}_{m}$, thus forming the noise-corrupted multimodal input $U^{1}_{m}\in\mathbb{R}^{T_{m}\times d_{m}}$.
@@ 3.3 Dominant Modality based Multimodal Learning
Inspired by previous work ALMT, we hypothesize that model robustness improves when the integrity of the dominant modality is preserved despite varying noise levels. We improve ALMT based on the designed DMC module and Reconstructor, thus implementing dominant modality based DMML module for sentiment analysis under random data missing scenarios. Here, we mainly introduce the parts that differ from ALMT. Further details are available in Zhang et al. (2023).
@@
$H_{hyper}^{i}=H_{hyper}^{i-1}+\operatorname{MHA}(H_{d}^{i},H_{a}^{1})+\operatorname{MHA}(H_{d}^{i},H_{v}^{1}),$ (3)
where $\operatorname{MHA}(\cdot)$ represents multi-head attention, $H_{hyper}^{i}\in\mathbb{R}^{T\times d}$ is the hyper-modality feature. Note that the feature $H_{hyper}^{0}\in\mathbb{R}^{T\times d}$ is a random initialized vector.
@@ 3.4 Dominant Modality Correction
Completeness Check. We apply an encoder $\operatorname{E}_{cc}$ that consists of a Transformer encoder with a depth of two layers and a classifier for completeness check. For example, if the missing rate of dominate modality is 0.3, the label of completeness is 0.7. This completeness prediction $w$ can be obtained as follows:
@@ 3.6 Overall Learning Objectives
where $\alpha$, $\beta$, $\gamma$ and $\delta$ are hyperparameters. On MOSI and MOSEI datasets, we empirically set them to 0.9, 0.8, 0.1, and 1.0, respectively. On the SIMS dataset, we empirically set them to 0.9, 0.6, 0.1, and 1.0, respectively.
@@ 4 Experiments and Analysis
In this section, we provide a comprehensive and fair comparison between the proposed LNLN and previous representative MSA methods on MOSI (Zadeh et al., 2016), MOSEI (Zadeh et al., 2018) and SIMS (Yu et al., 2020) datasets.
@@ 4.1 Datasets
MOSEI. The dataset consists of 22,856 video clips sourced from YouTube. The sample is divided into 16,326 clips for training, 1,871 for validation, and 4,659 for testing. Each clip is labeled with a score, ranging from -3, denoting the strongly negative, to 3, denoting the strongly positive.
@@ 4.2 Evaluation Settings and Criteria
For evaluation criteria, we report the binary classification accuracy (Acc-2), the F1 score associated with Acc-2, and the mean absolute error (MAE). For Acc-2, we calculated accuracy and F1 in two ways: negative/positive (left-side value of /) and negative/non-negative (right-side value of /) on the MOSI and MOSEI datasets, respectively. Additionally, we provide the three-class accuracy (Acc-3), the seven-class accuracy (Acc-7), and the correlation of the model’s prediction with humans (Corr) on the MOSI dataset. For the SIMS dataset, we report Acc-3, the five-class accuracy (Acc-5), and Corr. Due to the distinct focus of regression and classification metrics on different aspects of model performance, a model achieving the lowest error on regression metrics may not necessarily exhibit optimal performance on classification metrics. To comprehensively reflect the model capabilities, we select the best-performing checkpoint for each type of metric across all models in the comparisons, thus capturing the peak performance of both regression and classification aspects independently.
4.3 Implementation Details